How do i write the equation of a line that passes through (2,-9) and is perpendicular to the line (2x-9y=18)
Answer by scott8148(6628)

perpendicular means you can swap the x and y coefficients and change one of the signs

9x + 2y = ?

to find the new constant term, plug in the given point ___ 9(2) + 2(-9) = ? ___ 0 = ?

9x + 2y = 0

or, for those not comfortable with zero, ___ 2y = -9x ___ y = (-9/2)x